# AdaSteer: Your Aligned LLM is Inherently an Adaptive Jailbreak Defender

### v1

@@ Abstract
Despite extensive efforts in safety alignment, large language models (LLMs) remain vulnerable to jailbreak attacks. Activation steering offers a training-free defense method but relies on fixed steering coefficients, resulting in suboptimal protection and increased false rejections of benign inputs. To address this, we propose AdaSteer, an adaptive activation steering method that dynamically adjusts model behavior based on input characteristics. We identify two key properties: Rejection Law (R-Law), which shows that stronger steering is needed for jailbreak inputs opposing the rejection direction, and Harmfulness Law (H-Law), which differentiates adversarial and benign inputs. AdaSteer steers input representations along both the Rejection Direction (RD) and Harmfulness Direction (HD), with adaptive coefficients learned via logistic regression, ensuring robust jailbreak defense while preserving benign input handling. Experiments on LLaMA-3.1, Gemma-2, and Qwen2.5 show that AdaSteer outperforms baseline methods across multiple jailbreak attacks with minimal impact on utility. Our results highlight the potential of interpretable model internals for real-time, flexible safety enforcement in LLMs. WARNING: This paper may contain content that is offensive and harmful.
@@
Weixiang Zhao1thanks:    Equal contribution, Jiahe Guo1footnotemark: , Yulin Hu1, Yang Deng2, An Zhang3, Xingyu Sui1 Xinyang Han1, Yanyan Zhao1, Bing Qin1, Tat-Seng Chua3, Ting Liu1 1Harbin Institute of Technology, 2Singapore Management University 3National University of Singapore {wxzhao, jhguo, yyzhao}@ir.hit.edu.cn
@@ 1 Introduction
To address the first challenge, we conduct an empirical analysis of different types of jailbreak inputs along the RD within three safety-aligned LLMs: LLaMA-3.1 (Dubey et al., 2024), Gemma-2 (Team et al., 2024), and Qwen2.5 (Yang et al., 2024). As shown in Figure 1(b), we identity RD using contrastive pairs of complied (red cluster) and rejected (yellow cluster) harmful instructions via the difference-in-means technique (Belrose, 2023). We surprisingly find that different jailbreak types exhibit distinct patterns along RD, which can be summarizd as the Rejection Law (R-Law): {mdframed}[backgroundcolor=mycolor_blue!50, hidealllines=true] Rejection Law: Along RD, jailbreak types that are positioned further against the rejection direction are more difficult for the backbone model to defend against. Thus, R-Law can be leveraged as: the farther an input is along RD against the rejection direction, (i.e., the more adversary it is), the stronger rejection steering should be applied to enforce rejection.
However, solely depending on R-Law can not solve the second challenge as benign inputs can sometimes also exhibit distributions that oppose the rejection direction along RD, making them appear similar to jailbreak inputs. This directly motivates us to identity and leverage HD, reflecting the harmfulness of different inputs accordingly. Similarly, we obtain HD by contrasting complied harmful instructions with benign ones (blue cluster) and Harmfulness Law (H-Law) is derived: {mdframed}[backgroundcolor=mycolor_orange!50, hidealllines=true] Harmfulness Law: Along HD, jailbreak inputs shift further toward harmfulness compared to benign inputs (blue cluster), confirming their harmful nature and distinguishing them from benign queries. Since HD represents the backbone’s compliance behavior—identified by benign and harmful inputs that are both complied by the model—H-Law can be interpreted and leveraged as follows: the farther an input is along HD against the harmfulness direction, (i.e., the safer it is), the stronger the compliance steering should be applied along HD.
@@
Experiments on LLaMA-3.1-8B-Instruct (Dubey et al., 2024), Gemma-2-9B-it (Team et al., 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024) validate that R-Law and A-Law hold broadly. AdaSteer consistently outperforms baseline methods in jailbreak defense across 7 attack strategies. Furthermore, AdaSteer minimally affects the model’s performance on benign inputs, ensuring its utility remains intact. Our work serves as a concrete demonstration that insights gained from interpreting model internals can have practical applications and well-aligned LLMs hold significant potential to function as adaptive jailbreak defenders.
@@ 2 Preliminaries
A jailbreak attack seeks to craft an adversarial prompt $s^{\prime}=\mathcal{A}(s_{0})$, where $\mathcal{A}$ represents an attack method and $s_{0}$ is a vanilla harmful prompt. The objective is to induce the LLM to to generate a harmful response that aligns with the malicious intent of $s_{0}$, bypassing built-in safety mechanisms. Conversely, a jailbreak defense aims to protect the model against such adversarial manipulations.
@@ 3.1 Overview
$\boldsymbol{h\textquoteright}_{i}^{l}=\boldsymbol{h}^{l}_{i}+\lambda_{r}\boldsymbol{v}^{l}_{\text{RD}}+\lambda_{c}\boldsymbol{v}^{l}_{\text{HD}}$ (1)
@@ 3.2 Rejection Direction (RD), $\boldsymbol{v}_{\text{RD}}$ and $\lambda_{r}$
{mdframed}
[backgroundcolor=mycolor_blue!50, hidealllines=true] Rejection Law: Inputs that are positioned further in the negative direction against RD require a greater rejection steering coefficient $\lambda_{r}$ to induce rejection behavior.
@@ 3.3 Harmfulness Direction (HD), $\boldsymbol{v}_{\text{HD}}$ and $\lambda_{c}$
As shown in Figure 2, along the HD direction (x-axis), we also define the Harmful Compliance Center (red point) as the origin. The leftward direction represents less harmful (positive), while the rightward direction represents increased harmfulness (negative). Each input is projected onto the HD, yielding a coordinate $pos_{\text{HD}}$. On HD, we notice that jailbreak inputs generally have smaller $pos_{\text{HD}}$ values, whereas benign inputs, tend to have larger $dis_{\text{HD}}$ values, which can be summarized as the following Harmfulness Law. {mdframed}[backgroundcolor=mycolor_orange!50, hidealllines=true] Harmfulness Law: Inputs that are positioned further in the positive direction along HD require a greater compliance steering coefficient $\lambda_{c}$ to encourage compliance.
@@ 4.1 Experimental Setup
LLaMA-3.1 57 30 0 60 61 67 37 38.14 94.40 50.00
ROSE 100 83 51 94 85 61 85 79.86 90.45 2.80
Self-CD 94 67 5 66 67 43 43 55.00 93.75 2.30
Jailbreak Antidote 92 100 61 94 79 44 66 76.57 91.45 45.90
Surgical 100 75 10 88 84 82 91 75.71 82.35 47.30
InferAligner 85 90 0 92 77 82 77 71.86 80.45 47.20
CAST 100 100 0 66 76 46 56 63.43 95.00 37.70
AdaSteer (Ours) 100 100 82 90 85 100 86 91.86 97.85 50.01
@@
ROSE 99 52 8 86 58 12 0 45.00 97.00 1.00
Self-CD 69 50 2 82 54 6 0 37.57 96.00 1.00
Jailbreak Antidote 88 86 72 100 60 78 3 69.57 92.15 42.90
Surgical 94 41 0 82 47 13 3 40.00 95.25 47.90
InferAligner 100 98 0 98 60 94 11 65.86 93.40 48.80
@@
Gemma-2 6 31 0 90 57 1 27 30.29 86.25 50.00
ROSE 7 50 25 100 67 20 87 50.86 81.75 2.00
Self-CD 4 25 0 90 56 0 46 31.57 85.25 1.80
Jailbreak Antidote 6 47 0 98 61 1 78 41.57 83.35 47.30
Surgical 99 100 14 98 68 96 78 79.00 90.45 39.00
InferAligner 31 100 24 100 85 93 62 70.71 74.45 48.50
CAST 8 35 0 94 65 4 33 34.14 81.95 50.30
@@
$d_{\textbf{RD}}$ $pos_{\text{RD}}$71.77 -74.84 -72.16 -26.36 -63.80 -68.85 -65.07 -40.65 -45.62 -50.96
$\lambda_{r}$0.21 0.22 0.20 0.08 0.14 0.17 0.13 0.08 0.08 0.09
$d_{\textbf{HD}}$ $pos_{\text{HD}}$17.51 -17.36 -12.78 -17.01 -15.36 -14.74 -25.55 18.36 15.04 5.98
@@ 4.3 Analysis of Adaptive Steering
AdaSteer 91.86 97.80 50.01
w/o $\boldsymbol{v}_{\text{RD}}$ 39.57 98.55 50.70
w/o $\boldsymbol{v}_{\text{HD}}$ 91.57 74.35 45.72
@@
w/ reverse $\boldsymbol{v}_{\text{RD}}$ 87.43 90.55 48.05
@@
AdaSteer 85.86 92.75 48.28
w/o $\boldsymbol{v}_{\text{RD}}$ 56.57 88.65 49.99
w/o $\boldsymbol{v}_{\text{HD}}$ 92.14 90.15 33.08
@@ 4.4 Steering Vector Analysis
The results show that removing $\boldsymbol{v}_{\text{RD}}$ lowers jailbreak resistance, confirming its role in reinforcing rejection behavior. Conversely, removing $\boldsymbol{v}_{\text{HD}}$ significantly degrades utility, indicating that compliance steering along HD is crucial for reducing false rejections. The reverse $\boldsymbol{v}_{\text{RD}}$ setting achieves comparable jailbreak defense but sacrifices utility, demonstrating that simply inverting the rejection vector is suboptimal for distinguishing benign inputs. These findings validate the necessity of steering along both rejection and harmfulness direction for achieving robust and adaptive jailbreak defense.
@@ 4.6 Inference Efficiency Analysis
To evaluate the efficiency of different jailbreak defense methods, we compare their total inference time relative to the original model while maintaining a uniform batch size of 64 for methods that support batch inference. For methods that do not support batching, we use the ratio of their total inference time to that of the original model as an efficiency indicator. The trade-off between inference efficiency and jailbreak defense success rate (DSR) is visualized in Figure 4. AdaSteer is positioned in the upper-left region of the plot, demonstrating that it achieves a strong balance between safety and efficiency. Unlike other high-performing defenses that introduce significant computational overhead, AdaSteer maintains high DSR without excessive inference cost. This highlights its practicality as a scalable and efficient solution for enhancing model security in real-world deployments.
@@ 5 Related Works
Steering representation within LLMs has garnered increasing attention due to its transparency and lightweight properties (Zou et al., 2023a). Exist works mainly adopt static steering with a fixed coefficient exerted on the extracted refusal vectors for jailbreak defense (Zheng et al., 2024; Qian et al., 2024; Stickland et al., 2024; Li et al., 2025; Shen et al., 2025). Although few works explore more fine-grained steering control, they are still narrowed within vanilla harmful prompt scenario (Bhattacharjee et al., 2024; Wang et al., 2024c; Lee et al., 2025), leaving the more challenging jailbreak attacks under-explored.
@@ 6 Conclusion
In this work, we propose AdaSteer, a dual-direction adaptive activation steering method that enhances jailbreak defense in LLMs while maintaining their utility. By identifying two key properties—Rejection Law and Harmfulness Law—we show that jailbreak inputs exhibit distinct behaviors in activation space, allowing for dynamic, input-aware steering along the Rejection Direction and Harmfulness Direction. Extensive experiments on LLaMA-3.1, Gemma-2, and Qwen2.5 confirm that AdaSteer consistently outperforms baseline defenses across diverse jailbreak strategies, demonstrating its effectiveness and scalability.
@@ A.2 Benchmarks
  • Cipher (Yuan et al., 2024) Cipher is a jailbreak technique that leverages vulnerabilities in large language models (LLMs) by employing encoding methods to circumvent content filters and safety protocols. This approach embeds encoded or obfuscated commands within prompts, enabling them to slip past detection systems. In our experiments, we utilize EasyJailbreak along with 100 harmful queries from AdvBench to create the jailbreak inputs.
  • GCG (Zou et al., 2023b) GCG, which stands for Greedy Coordinate Gradient, is a method used to jailbreak LLMs. This approach automatically creates discrete adversarial tokens. During the optimization process, it selects the suffix that results in the lowest loss. Although it lost some readability, it achieved a good attack effect. In our experiments, we utilize EasyJailbreak along with 100 harmful queries from AdvBench to create the jailbreak inputs.
@@
  • OKTest (Röttger et al., 2024) It includes 300 test samples featuring safe questions that incorporate harmful and sensitive words.
@@ Appendix C Implementation Details
For the identified vector $v_{\text{HD}}$ undergoing the projection process, we evaluate its effect by introducing it with a positive coefficient $\lambda_{c}$ > 0. If this does not lead to an improvement in the model’s compliance, we reverse its direction.
Regarding the number of compiled harmful examples, we were surprised to find that even a small number of such examples was sufficient to achieve the desired effect. In our main experiments, we used 15 compiled harmful examples for LLaMA-3.1-8B-Instruct, 13 for Qwen2.5-7B-Instruct, and 5 for Gemma-2-9B-it. In addition, we included an equal number of rejected harmful examples and complied benign data for each model.
To find the value of $\lambda_{r}$ necessary for the model to reject all inputs in Figure 2, we first calculate the average position of harmful inputs rejected on the RD. This average position represents the precise location of the harmful rejection center. For logistic regression fitting, we performed a simple fit on RD using the mentioned compiled harmful examples. We then conducted a grid search on the Validation Set described in A.3. Similarly, after dynamically applying RD, we label $\lambda_{c}$ as the threshold at which benign inputs that were previously mistakenly rejected begin to be correctly accepted. We then fit the $\lambda_{c}$ curve and adjust it using grid search.
We want to emphasize that $\lambda_{r}$ and $\lambda_{c}$ should not be infinitely large or small, because once they reach a certain value, further increasing or decreasing them becomes meaningless and may even lead to decoding failure. To avoid this, we set upper and lower limit $\lambda_{r}$ and $\lambda_{c}$ values for truncation on the fitted logistic regression curve. Therefore, the average $pos$ and $\lambda$ in the Table 2, Table 7 and Table 8 might not exhibit strict linearity, but each input still receives the necessary steering force.

### v2

@@ Abstract
Despite extensive efforts in safety alignment, large language models (LLMs) remain vulnerable to jailbreak attacks. Activation steering offers a training-free defense method but relies on fixed steering coefficients, resulting in suboptimal protection and increased false rejections of benign inputs. To address this, we propose AdaSteer, an adaptive activation steering method that dynamically adjusts model behavior based on input characteristics. We identify two key properties: Rejection Law (R-Law), which shows that stronger steering is needed for jailbreak inputs opposing the rejection direction, and Harmfulness Law (H-Law), which differentiates adversarial and benign inputs. AdaSteer steers input representations along both the Rejection Direction (RD) and Harmfulness Direction (HD), with adaptive coefficients learned via logistic regression, ensuring robust jailbreak defense while preserving benign input handling. Experiments on LLaMA-3.1, Gemma-2, and Qwen2.5 show that AdaSteer outperforms baseline methods across multiple jailbreak attacks with minimal impact on utility. Our results highlight the potential of interpretable model internals for real-time, flexible safety enforcement in LLMs. Our code is available at: https://github.com/MuyuenLP/AdaSteer. WARNING: This paper may contain content that is offensive and harmful.
@@
Weixiang Zhao1thanks: Equal contribution, Jiahe Guo1footnotemark: , Yulin Hu1, Yang Deng2, An Zhang3, Xingyu Sui1 Xinyang Han1, Yanyan Zhao1thanks: Corresponding author, Bing Qin1, Tat-Seng Chua3, Ting Liu1 1Harbin Institute of Technology, 2Singapore Management University 3National University of Singapore {wxzhao, jhguo, yyzhao}@ir.hit.edu.cn
@@ 1 Introduction
To address the first challenge, we conduct an empirical analysis of different types of jailbreak inputs along the RD within three safety-aligned LLMs: LLaMA-3.1 (Dubey et al., 2024), Gemma-2 (Team et al., 2024), and Qwen2.5 (Yang et al., 2024). As shown in Figure 1(b), we identity RD using contrastive pairs of complied (red cluster) and rejected (yellow cluster) harmful instructions via the difference-in-means technique (Belrose, 2023). We surprisingly find that different jailbreak types exhibit distinct patterns along RD, which can be summarizd as the Rejection Law (R-Law): Rejection Law: Along RD, jailbreak types that are positioned further against the rejection direction are more difficult for the backbone model to defend against. Thus, R-Law can be leveraged as: the farther an input is along RD against the rejection direction, (i.e., the more adversary it is), the stronger rejection steering should be applied to enforce rejection.
However, solely depending on R-Law can not solve the second challenge as benign inputs can sometimes also exhibit distributions that oppose the rejection direction along RD, making them appear similar to jailbreak inputs. This directly motivates us to identity and leverage HD, reflecting the harmfulness of different inputs accordingly. Similarly, we obtain HD by contrasting complied harmful instructions with benign ones (blue cluster) and Harmfulness Law (H-Law) is derived: Harmfulness Law: Along HD, jailbreak inputs shift further toward harmfulness compared to benign inputs (blue cluster), confirming their harmful nature and distinguishing them from benign queries. Since HD represents the backbone’s compliance behavior—identified by benign and harmful inputs that are both complied by the model—H-Law can be interpreted and leveraged as follows: the farther an input is along HD against the harmfulness direction, (i.e., the safer it is), the stronger the compliance steering should be applied along HD.
@@
Experiments on LLaMA-3.1-8B-Instruct (Dubey et al., 2024), Gemma-2-9B-it (Team et al., 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024) validate that R-Law and H-Law hold broadly. AdaSteer consistently outperforms baseline methods in jailbreak defense across 7 attack strategies. Furthermore, AdaSteer minimally affects the model’s performance on benign inputs, ensuring its utility remains intact. Our work serves as a concrete demonstration that insights gained from interpreting model internals can have practical applications and well-aligned LLMs hold significant potential to function as adaptive jailbreak defenders.
@@ 2 Preliminaries
A jailbreak attack seeks to craft an adversarial prompt $s^{\prime}=\mathcal{A}(s_{0})$, where $\mathcal{A}$ represents an attack method and $s_{0}$ is a vanilla harmful prompt. The objective is to induce the LLM to generate a harmful response that aligns with the malicious intent of $s_{0}$, bypassing built-in safety mechanisms. Conversely, a jailbreak defense aims to protect the model against such adversarial manipulations.
@@ 3.1 Overview
$\boldsymbol{h’}_{i}^{l}=\boldsymbol{h}^{l}_{i}+\lambda_{r}\boldsymbol{v}^{l}_{\text{RD}}+\lambda_{c}\boldsymbol{v}^{l}_{\text{HD}}$ (1)
@@ 3.2 Rejection Direction (RD), $\boldsymbol{v}_{\text{RD}}$ and $\lambda_{r}$
Rejection Law: Inputs that are positioned further in the negative direction against RD require a greater rejection steering coefficient $\lambda_{r}$ to induce rejection behavior.
@@ 3.3 Harmfulness Direction (HD), $\boldsymbol{v}_{\text{HD}}$ and $\lambda_{c}$
As shown in Figure 2, along the HD direction (x-axis), we also define the Harmful Compliance Center (red point) as the origin. The leftward direction represents less harmful (positive), while the rightward direction represents increased harmfulness (negative). Each input is projected onto the HD, yielding a coordinate $pos_{\text{HD}}$. On HD, we notice that jailbreak inputs generally have smaller $pos_{\text{HD}}$ values, whereas benign inputs, tend to have larger $pos_{\text{HD}}$ values, which can be summarized as the following Harmfulness Law. Harmfulness Law: Inputs that are positioned further in the positive direction along HD require a greater compliance steering coefficient $\lambda_{c}$ to encourage compliance.
@@ 4.1 Experimental Setup
LLaMA-3.1 57 30 0 60 61 22 37 38.14 94.40 50.00
ROSE 100 83 51 94 85 61 85 79.86 90.47 2.81
Self-CD 94 67 5 66 67 43 43 55.00 93.74 2.27
Jailbreak Antidote 92 100 61 94 79 44 66 76.57 91.44 45.93
Surgical 100 75 10 88 84 82 91 75.71 82.37 47.29
InferAligner 85 90 0 92 77 82 77 71.86 80.47 47.19
CAST 100 100 0 66 76 46 56 63.43 95.00 37.76
AdaSteer (Ours) 100 100 82 90 85 100 86 91.86 97.87 50.01
@@
ROSE 99 52 8 86 58 12 0 45.00 97.00 1.03
Self-CD 69 50 2 82 54 6 0 37.57 96.00 0.96
Jailbreak Antidote 88 86 72 100 60 78 3 69.57 93.17 42.86
Surgical 94 41 0 82 47 13 3 40.00 95.24 48.85
InferAligner 100 98 0 98 60 94 11 65.86 93.40 48.43
@@
Gemma-2 6 31 0 90 57 1 27 30.29 86.27 50.00
ROSE 7 50 25 100 67 20 87 50.86 81.74 1.98
Self-CD 4 25 0 90 56 0 46 31.57 85.24 1.75
Jailbreak Antidote 6 47 0 98 61 1 78 41.57 83.34 47.33
Surgical 99 100 14 98 68 96 78 79.00 90.57 38.98
InferAligner 31 100 24 100 85 93 62 70.71 74.44 48.48
CAST 8 35 0 94 65 4 33 34.14 81.94 50.32
@@
$d_{\textbf{RD}}$ $pos_{\text{RD}}$ -71.77 -74.84 -72.16 -26.36 -63.80 -68.85 -65.07 -40.65 -45.62 -50.96
$\lambda_{r}$ -0.21 0.22 0.20 0.08 0.14 0.17 0.13 0.08 0.08 0.09
$d_{\textbf{HD}}$ $pos_{\text{HD}}$ -17.51 -17.36 -12.78 -17.01 -15.36 -14.74 -25.55 18.36 15.04 5.98
@@ 4.3 Analysis of Adaptive Steering
AdaSteer 91.86 97.87 50.01
w/o $\boldsymbol{v}_{\text{RD}}$ 39.57 98.54 50.70
w/o $\boldsymbol{v}_{\text{HD}}$ 91.57 74.37 45.72
@@
w/ reverse $\boldsymbol{v}_{\text{RD}}$ 87.43 90.54 48.05
@@
AdaSteer 85.86 92.80 48.28
w/o $\boldsymbol{v}_{\text{RD}}$ 56.57 88.67 49.99
w/o $\boldsymbol{v}_{\text{HD}}$ 92.14 90.17 33.08
@@ 4.4 Steering Vector Analysis
The results show that removing $\boldsymbol{v}_{\text{RD}}$ lowers jailbreak resistance, confirming its role in reinforcing rejection behavior. Conversely, removing $\boldsymbol{v}_{\text{HD}}$ significantly degrades utility, indicating that compliance steering is crucial for reducing false rejections. The reverse $\boldsymbol{v}_{\text{RD}}$ setting achieves comparable jailbreak defense but sacrifices utility, demonstrating that simply inverting the rejection vector is suboptimal for distinguishing benign inputs. These findings validate the necessity of steering along both rejection and harmfulness direction for achieving robust and adaptive jailbreak defense.
@@ 4.6 Inference Efficiency Analysis
To evaluate the efficiency of different jailbreak defense methods, we compare their tokens per second (token/s) relative to the original model. We conduct our experiments on a single NVIDIA Tesla A100 GPU. For methods that support batch inference, we set the batch size to 64. The trade-off between inference efficiency and jailbreak defense success rate (DSR) is visualized in Figure 4. AdaSteer is positioned in the upper-right region of the plot, demonstrating that it achieves a strong balance between safety and efficiency. Unlike other high-performing defenses that introduce significant computational overhead, AdaSteer maintains high DSR without excessive inference cost, preserving a runtime speed close to that of the original model. This highlights its practicality as a scalable and efficient solution for enhancing model security in real-world deployments.
@@ 5 Related Works
Steering representation within LLMs has garnered increasing attention due to its transparency and lightweight properties (Zou et al., 2023a). This technique is grounded in the theoretical premise that LLMs encode features or concepts as linear directions in activation space (Mikolov et al., 2013; Park et al., 2024). Exist works mainly adopt static steering with a fixed coefficient exerted on the extracted refusal vectors for jailbreak defense (Zheng et al., 2024; Qian et al., 2024; Stickland et al., 2024; Li et al., 2025; Shen et al., 2025). Although few works explore more fine-grained steering control, they are still narrowed within vanilla harmful prompt scenario (Bhattacharjee et al., 2024; Wang et al., 2024c; Lee et al., 2025), leaving the more challenging jailbreak attacks under-explored.
@@ 6 Conclusion
In this work, we propose AdaSteer, a dual-direction adaptive activation steering method that enhances jailbreak defense in LLMs while maintaining their utility. By identifying two key properties—Rejection Law and Harmfulness Law—we show that jailbreak inputs exhibit distinct behaviors in activation space, allowing for dynamic, input-aware steering along the Rejection and Harmfulness Direction. Extensive experiments on LLaMA-3.1, Gemma-2, and Qwen2.5 confirm that AdaSteer outperforms baselines across diverse jailbreak strategies, demonstrating its effectiveness and scalability.
@@
Acknowledgments
We thank the anonymous reviewers for their comments and suggestions. This work was supported by the New Generation Artificial Intelligence-National Science and Technology Major Project 2023ZD0121100, the National Natural Science Foundation of China (NSFC) via grant 62441614 and 62176078, the Fundamental Research Funds for the Central Universities, and the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (No. MSS24C012).
@@ A.2 Benchmarks
  • Cipher (Yuan et al., 2024) Cipher is a jailbreak technique that leverages vulnerabilities in large language models (LLMs) by employing encoding methods to circumvent content filters and safety protocols. This approach embeds encoded or obfuscated commands within prompts, enabling them to slip past detection systems. In our experiments, we utilize EasyJailbreak along with 25 harmful queries from AdvBench to create the jailbreak inputs.
  • GCG (Zou et al., 2023b) GCG, which stands for Greedy Coordinate Gradient, is a method used to jailbreak LLMs. This approach automatically creates discrete adversarial tokens. During the optimization process, it selects the suffix that results in the lowest loss. Although it lost some readability, it achieved a good attack effect. In our experiments, we utilize EasyJailbreak along with 50 harmful queries from AdvBench to create the jailbreak inputs.
@@
  • OKTest (Shi et al., 2024b) It includes 300 test samples featuring safe questions that incorporate harmful and sensitive words.
@@ Appendix C Implementation Details
To determine the value of $\lambda_{r}$ required for the model to reject all jailbreak inputs in Figure 2, we first categorize the harmful inputs into those that are rejected and those that are complied with. We then calculate the average position of the rejected harmful inputs on the RD. This average position represents the exact location of the harmful rejection center. Next, we determine the $\lambda_{r}$ needed to draw the complied jailbreak inputs and toward this center. Similarly, we label the harmful inputs that were complied with in the same manner, for use in subsequent logistic regression fitting.
For logistic regression fitting, we performed a simple fit on RD using the mentioned compiled harmful examples. Regarding the number of compiled harmful examples, we are surprised to find that even a small number of such examples is sufficient to achieve the desired effect. In our main experiments, we use 15 compiled harmful examples for LLaMA-3.1-8B-Instruct, 13 for Qwen2.5-7B-Instruct, and 5 for Gemma-2-9B-it. We then conducted a grid search on the Validation Set described in A.3. Similarly, after dynamically applying RD, we label $\lambda_{c}$ as the threshold at which benign inputs that were previously mistakenly rejected begin to be correctly accepted — for example, 158 such cases for LLaMA-3.1-8B-Instruct. We then fit the $\lambda_{c}$ curve and adjust it using grid search.
We want to emphasize that $\lambda_{r}$ and $\lambda_{c}$ should not be infinitely large or small, because once they reach a certain value, further increasing or decreasing them becomes meaningless and may even lead to decoding failure. To avoid this, we set upper and lower limit $\lambda_{r}$ and $\lambda_{c}$ values for truncation on the fitted logistic regression curve. Therefore, the average pospos and $\lambda$ in the Table 2, Table 8 and Table 9 might not exhibit strict linearity, but each input still receives the necessary steering force.